# Predictability of viral load kinetics in the early phases of SARS-CoV-2 through a model-based approach

### v1

@@ Abstract
A pipeline to evaluate the evolution of viral dynamics based on a new model-driven approach has been developed in the present study. The proposed methods exploit real data and the multiscale structure of the infection dynamics to provide robust predictions of the epidemic dynamics. We focus on viral load kinetics whose dynamical features are typically available in the symptomatic stage of the infection. Hence, the epidemiological evolution is obtained by relying on a compartmental approach characterized by a varying infection rate to estimate early-stage viral load dynamics, of which few data are available. We test the proposed approach with real data of SARS-CoV-2 viral load kinetics collected from patients living in an Italian province. The considered database refers to early-phase infections, whose viral load kinetics are not affected by mass vaccination policies in Italy. Our contribution is devoted to provide an effective computational pipeline to evaluate in real time the evolution of infectivity. Comprehending the factors influencing the in-host viral dynamics represents a fundamental tool to provide robust public health strategies. This pilot study could be implemented in further investigations involving other respiratory viruses, to better clarify the process of viral dynamics as a preparatory action for future pandemics.
@@ 1 Introduction
Infection dynamics are typically shaped by heterogeneous and interconnected factors, including exogenous individual-based aspects such as the individual response to threat, the collective compliance to non-pharmaceutical interventions, and variable endogenous characteristics of a pathogen [6, 7, 8, 12, 21, 22, 26, 37, 44]. Indeed, the infectivity of individuals may vary strongly from the onset of the infection to its end. In a purely data-oriented approach, the day of first contact with an infected individual is typically unknown or unreported and the infectivity levels may vary in time. Therefore, the trend of an epidemic naturally involves a plurality of timescales: one that is linked to the evolution of infected cases among the population and a second related with the change of infectivity related to in-host viral dynamics [25, 33, 4].
@@
In this direction, viral load (VL) dynamics represent a significant aspect for the disease progression and transmission. The evolution of VL represents a quantitative marker for assessing viral kinetics [20, 36, 40]. Nevertheless, the reconstruction of these trajectories is a challenging problem since data are incomplete and biased toward the peak viral load, that is typically related to the onset of symptoms in infected patients. We mention in this direction the empirical studies, see e.g. [23]. Hence, VL kinetics are strongly affected by uncertainties stemming from the data assimilation processes. Existing methods are not well-suited to consider possible mutations of the virus, resulting in a modification of the infectiousness curve associated with each individual illness. Therefore, a purely data-oriented approach should be complemented by a multiscale model-oriented approach to provide robust long-term predictions of epidemiological dynamics [28]. In the last few years, several mathematical approaches have been designed to assess the impact of infectivity dynamics at the population level [14, 32]. These works are based on kinetic-type equations and on simplified dynamics for which observable quantities are analytically computable. The developed methods are capable to link agent-based dynamics to available data, providing effective transition rates in macroscopic compartmental models, see also [3, 5, 17, 19, 43, 44].
@@
In this study, we focus on the definition of a new pipeline capable of extracting information starting from viral load kinetics through a model-driven approach that couples the multiple timescales of infection dynamics. In detail, we exploit VL kinetics obtained in a series of SARS-CoV-2 infections that are not affected by exogenous influencing factors in the dynamics, like e.g. the implementation of mass vaccination policies. In detail, we interface data-oriented viral load dynamics with a simple compartmental model with age of infection [15, 42] to describe the evolution of infected cases from SARS-CoV-2 in Italy. A new approach to evaluate the dynamics of the viral load has been developed and compared with original data obtained during the initial waves of the SARS-CoV-2 pandemic.
2 Materials and methods
@@ 2.1 Samples and SARS-CoV-2 kinetics
Since February 2020, Fondazione IRCCS Policlinico San Matteo (Pavia, Italy) has been identified as reference regional center for the diagnosis of SARS-CoV-2 infections, through the analysis of respiratory samples. In these analyses, RNA was extracted by using the MGISP-960 automated workstation and the MGI Easy Magnetic Beads Virus DNA/RNA Extraction Kit (MGI Technologies, Shenzhen, China). SARS-CoV-2 RNA was detected using the SARS-CoV-2 variants ELITe MGB kit (ELITechGroup, Puteaux, France; cat. no. RTS170ING) targeting ORF8 and RdRp gene. Reactions were carried out on the CFX96 Touch Real-time PCR detection system (BioRad, Mississauga, ON, Canada).
@@
$1\ \frac{\textrm{copies}}{\textrm{ml}}=16.6\times 10^{-0.2814\ \times\ 1\textrm{Ct}\ +\ 11.232},$
@@ 2.2 Statistical analysis
Numerical variables are described as mean $\pm$ standard deviation, categorical variables as count and percentages. The comparison of the two periods of time considered is performed in terms of t-test for quantitative variables and Fisher exact test for quantitative variables. Considering the longitudinal structure of the data, they are presented as daily averaged values. The comparison between the two periods and the investigation of the influence of age and gender on the kinetics are performed in terms of a linear mixed-effects model. The temporal evolution of VL kinetics is obtained as a Gamma distribution using the Matlab built-in fit function, using the LAR method to ensure that the extreme values have a lesser influence on the result. The significance threshold is set as $p<0.05$.
3 Results
@@ 3.1 Data description
The study analyzed a total of 233 viral load kinetics, corresponding to approximately 700 respiratory samples, collected from patients across two distinct periods. The first period goes from November to December 2020, before the introduction of variants of concern (VOCs), and included 71 patients. The second period, from January to May 2021, during the circulation of VOCs, involved 162 patients.
@@
[…]
3.2 Data-oriented viral load kinetics
To understand what the proper target shape for the theoretical function $\beta$ could be (and its evolution over the two observation periods), we initially infer it from the best fit of the individuals’ average viral load, from each one of the considered temporal windows. Based on existing works in the medical literature about the shape of the SARS-CoV-2 infectiousness [11, 37], we expect $\beta$ to be characterized by a rapid growth in the few days after the first contact with an infective individual, until the infection reaches its maximum, followed by a slower decay, up to the complete negativization of the subject after several days. The issue with the raw data is that we cannot be sure that the first molecular test of each patient corresponds to their infections’s peak. Therefore, we loose any deterministic information on both the day in which the infection reaches its maximum and the value of this maximum. To rule out part of this randomness, we reorder the infection cycle of each subject based on their first negative test, rather than their first positive one. Indeed, while it remains unclear when the actual negativization happened, we know for certain that the value of the viral load on that day must be 0. In Figure 1 we present the raw data for the viral load kinetics of each individual patient ordered by the first negativization, over the two distinct periods of time during which their infectiousness was monitored.
Figure 1: Evolution of individuals’ viral load kinetics over time, with an ordering based on the first negative test taken by the infective patients. We distinguish between the two observation periods of November–December 2020, accounting for 71 patients (left), and January–May 2021, accounting for 162 patients (right).
Starting from this reordering, we consider the daily average of the viral load with respect to the number of patients that got tested on that day. We then infer the shape of $\beta$ as the best fit of these averages, weighted by the daily patients’ count. This means that the more patients got tested on one specific day, the more importance the corresponding average viral load value will have in the fitting. The outcomes of this analysis are shown in Figure 2. On average, the infection persists for 25 days at most and its peak is reached 3 to 4 days after the first contact with an infected individual (4 days for the viral loads considered from the period November–December 2020, and 3 days for those belonging to the window January–May 2021). The red curves represent the best weighted fits for the daily averaged viral loads and have the shape of a Gamma distribution
@@ 3.2 Data-oriented viral load kinetics
where $\alpha_{0}=10^{7}$ is a scale factor used to capture the magnitude of individuals’ viral load and the coefficients $\alpha_{1}\in\mathbb{R}$, $\alpha_{2}$, $\alpha_{3}\geq 0$ are reported in Table 1.
@@
Table 1: Optimised coefficients for viral load kinetics, for the two observation periods of November–December 2020 (left) and January–May 2021 (right).
@@ 3.3 Epidemic model with uncertain quantities
Compartmental models are classically defined to describe mathematically the spread of epidemics in a population [16]. The main idea of such models is to divide the population into different subgroups, each identified by a specific epidemiologically relevant status. Suitable transition rates are introduced to describe the switches between compartments. Generalizations of the mentioned compartmental models include additional factors characterizing the transmission dynamics, such as external influence, age structure, variable contact dynamics and mobility of agents, see e.g. [1, 2, 9, 18]. In the following, we will consider a SIR-type model with age of infection where the infectivity keeps track of the disease’s variability inside each individual and evolves in parallel with the main timescale of the epidemic. This allows to introduce a local incidence rate function of the age of infection, modelling the average viral load of the infected individuals and whose shape can be inferred by the information coming from real data on molecular tests. To take into account the structural uncertainties stemming from data-assimilation processes of VL kinetics, we extend the model approach to consider the presence of uncertain quantities as a structural feature of the epidemic dynamics. To date, few results are available regarding the development of uncertainty quantification (UQ) methods in epidemic systems, see [1, 10, 13, 38]. The development of UQ methods are based on an increased dimensionality of the problem which is made dependent on a random variable $\mathbf{z}\in\mathbb{R}^{d_{z}}$ of which the distribution $p(\mathbf{z})$ is known [24, 34, 41]. The extrapolation of statistics is classically obtained by looking at quantities of interest, with respect to $\mathbf{z}$, representing the expected solution of the problem.
At a given time $t\geq 0$, we subdivide the population into susceptible $S$, recovered $R$ and infected with variable infectiousness $i$. We then consider the new triple $(S(t,\mathbf{z}),i(t,\tau,\mathbf{z}),R(t,\mathbf{z}))$ of unknowns, solution to
@@
where now $\beta(\tau,\mathbf{z})$ and $\gamma(\tau,\mathbf{z})$ represent the uncertain infectiousness and recovery of the infected individuals, depending on the age of infection $\tau\geq 0$, while the constants $\bar{\beta}>0$ and $\bar{\gamma}>0$ enclose the average contact rate and duration of the illness. We observe that the observable number of infected also varies based on the new timescale $\tau\geq 0$, meaning that the quantity $i(t,\tau,\mathbf{z})\geq 0$ depends on both the timescale, and it is related to the total number of infected $I$ through the relation
@@
This model encapsulates the timescale of the epidemic with the one characterizing the evolution of the infection, in terms of infectiousness. For each $\mathbf{z}$, the system is finally completed with suitable initial and boundary conditions for the variables $S(t,\mathbf{z})$, $i(t,\tau,\mathbf{z})$ and $R(t,\mathbf{z})$. A natural choice is
@@
where $p(\mathbf{z})$ is the distribution of the random variable $\mathbf{z}$.
3.4 Inference of viral load dynamics based on uncertain data
@@
In this section we define a pipeline to effectively estimate actual viral load kinetics $\beta(\tau,\mathbf{z})\geq 0$, by using available epidemiological data and introducing uncertainty in the day of first contact, expressed by the coefficient $\alpha_{1}$ in (1). To this end, we consider an uncertain parameter $\alpha_{1}(\mathbf{z})\geq 0$ to make the local incidence rate random. Moreover, we take into account data from the COVID-19 pandemic, both at local and national level, by considering the evolution of infected in the province of Pavia and then extending our analysis to the whole territory of Italy, where the central government implemented several non-pharmaceutical interventions (NPIs) during the pandemic. In Table 2 we report the timeline of the main NPIs taken before the mass vaccination campaign started in mid 2021. It is however worth mentioning that local additional measures, like isolation of small portion of a territory, had often been implemented as well.
@@
Before advancing to the estimation of VL kinetics, it is crucial to address some key points. We stress the inherent complexity in calibrating epidemiological models, which is particularly demanding from a numerical perspective. Indeed, the process of aligning these models with real-world scenarios requires a careful approach to parameter estimation. Furthermore, available data on viral load kinetics tend to hide the initial evolution of this dynamics, since agents are typically tested after the manifestation of the symptoms. This bias is primarily attributed to limited testing capacity, resulting in a dataset that often reflects a conservative estimate. Consequently, any attempts to estimate the viral loads should be approached with a mindful consideration of these inherent limitations, coupled with a commitment to refining the model’s accuracy through continuous reassessment and adaptation.
Implemented NPI Date Average recovery rate
First lockdown March 9, 2020
Curfew measures October 22, 2020 14 days
Strengthening of measures March 10, 2021
@@
Table 2: Timeline of the main NPIs employed in Italy before the mass vaccination campaign started in mid 2021.
@@
We shall focus on three different periods of time. We start by analyzing the first COVID-19 epidemic wave in Italy, which began in February–March 2020. We then proceed with the study of the second wave of October 2020, and conclude with the third major wave of February–March 2021. In particular, we point out that our epidemiological data for the province of Pavia were only acquired up to January 18 2021, and we will therefore narrow down our analysis on the third wave to the sole case of Italy, for which we shall instead take advantage of the open access data collected by the Johns Hopkins University (github.com/CSSEGISandData/COVID-19), reporting in details the daily number of infected, recovered and deceased individuals in Italy up to August 4 2021. Let $\mathcal{C}_{z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{M}\}$ be a sample of the random variable $\mathbf{z}$ with probability density function $p(\mathbf{z})$. We have adopted a multi-level approach for estimating the parameters $(\alpha_{i})_{i=1,2,3}$ that characterize $\beta(\tau,\mathbf{z}_{j})\geq 0$, which is articulated in two steps.
@@
Step 1 – Determination of the epidemiological parameters. For starting, we initially assess the (deterministic) value of $\alpha_{1}$ dictated by the epidemic during the two weeks preceding any of the three main NPIs presented in Table 2, and we estimate at the same time the relevant epidemiological parameter $\bar{\beta}\geq 0$ which defines the baseline transmission dynamics, assuming no restrictions on social contacts were in place. Notice that we have to discard the variable $\mathbf{z}$ here, because we first need to ensure that the day of first contact $\alpha_{1}$ is compatible with the epidemiological data, in order to quantify the shift between the infection and the epidemic timescales. For this, we look for the optimal $\alpha_{1}$ and $\bar{\beta}$ minimizing the relative weighted $L^{2}$ norm between the reported number of infected and recovered individuals (in Pavia or in Italy, depending on the local or national setting under consideration) which we denote $\hat{I}(t)$ and $\hat{R}(t)$, and the theoretical evolution of these quantities, $I(t)=\int_{\mathbb{R}_{+}}i(\tau)\mathrm{d}\tau$ and $R(t)$, given by the model (2) in absence of uncertainty. More precisely, we solve the following minimization problem
@@
over the time horizon $[t_{0},t_{f}]$ representing any of the three periods that can be identified with the initial spreading phase of the three epidemic waves under study: February 24–March 9 2020, October 7–October 22 2020, and February 24–March 10 2020. In particular, we choose here the infectiousness $\beta(\tau)$ to be the Gamma functions (1), obtained in the previous section from the observed viral loads. We use the fitting from November–December 2020 as initial guess for the optimization (5) during the first two epidemic waves, while we opt for the January–May 2021’s fitting to study the third wave. Notice that at this stage the parameters $\alpha_{2}$ and $\alpha_{3}$ are fixed from these two fits, since we only need to ensure a proper initial calibration of the time shift $\alpha_{1}$, to correctly align our model with the epidemiological data. Finally, we always consider a constant recovery rate $\bar{\gamma}=1/14$ to lower the complexity of the minimization. This value is consistent with the existing literature on the COVID-19 infection [12, 30, 44], for which the time to viral clearance has been reported to span approximately between 10 and 21 days on average over the different variants. In Table 3, we collect all the parameters obtained from this initial analysis, for each of the three waves.
@@
$\bar{\beta}$ 0.1486 0.1603
$\bar{\gamma}$ 1/14 1/14
@@
$\bar{\beta}$ 0.3992 0.1225 0.1323
$\bar{\gamma}$ 1/14 1/14 1/14
@@
Table 3: Epidemiological parameters and coefficients of the infectiousness fuction, for the three periods of time under consideration, both at a local (province of Pavia) and a national level (Italy).

### v2

# Predictability of viral load dynamics in the early phases of SARS-CoV-2 through a model-based approach

@@ Abstract
A pipeline to evaluate the evolution of viral dynamics based on a new model-driven approach has been developed in the present study. The proposed methods exploit real data and the multiscale structure of the infection dynamics to provide robust predictions of the epidemic dynamics. We focus on viral load kinetics whose dynamical features are typically available in the symptomatic stage of the infection. Hence, the epidemiological evolution is obtained by relying on a compartmental approach characterized by a varying infection rate to estimate early-stage viral load dynamics, of which few data are available. We test the proposed approach with real data of SARS-CoV-2 viral load kinetics collected from patients living in an Italian province. The considered database refers to early-phase infections, whose viral load kinetics have not been affected by the mass vaccination policies in Italy.
@@ 1 Introduction
Infection dynamics are typically shaped by heterogeneous and interconnected factors, including exogenous individual-based aspects such as the individual response to threat, the collective compliance to non-pharmaceutical interventions, and variable endogenous characteristics of a pathogen [6, 8, 9, 10, 14, 23, 24, 29, 44, 51]. Indeed, the infectivity of individuals may vary strongly from the onset of the infection to its end. In a purely data-oriented approach the day of first contact with an infected individual is typically unknown or unreported and the infectivity levels may vary in time. Therefore, the trend of an epidemic naturally involves a plurality of timescales: one that is linked to the evolution of infected cases among the population and a second related with the change of infectivity related to in-host viral dynamics [28, 36, 5].
@@
In this direction, viral load (VL) dynamics represent a significant aspect influencing the disease progression and transmission. The evolution of VL constitutes a quantitative marker for assessing viral kinetics [22, 43, 47]. Nevertheless, the reconstruction of these trajectories is a challenging problem since data are incomplete and biased toward the peak viral load, which is typically related to the onset of symptoms in infected patients. We mention in this direction the empirical studies, see e.g. [26]. Hence, VL kinetics are strongly affected by uncertainties stemming from the data assimilation processes. Existing methods are not well-suited to consider possible mutations of the virus, resulting in a modification of the infectiousness curve associated with each individual illness. Therefore, a purely data-oriented approach should be complemented by a multiscale model-oriented approach to provide robust long-term predictions of epidemiological dynamics [31]. In the last few years, several mathematical approaches have been designed to assess the impact of infectivity dynamics at the population level [16, 35]. These works are based on kinetic-type equations and on simplified dynamics for which observable quantities are analytically computable. The developed methods are capable to link agent-based dynamics to available data, providing effective transition rates in macroscopic compartmental models, see also [3, 4, 7, 19, 21, 50, 51].
@@
In this study, we focus on the definition of a new pipeline capable of extracting information starting from VL kinetics through a model-driven approach that couples the multiple timescales of infection dynamics. In detail, we exploit VL kinetics obtained in a series of SARS-CoV-2 infections that are not affected by exogenous influencing factors in the dynamics, like e.g. the implementation of mass vaccination policies. We interface data-oriented VL dynamics with a simple compartmental model involving age of infection [17, 49] to describe the evolution of infected cases from SARS-CoV-2 in Italy. A new approach to evaluate the dynamics of the viral load has been developed and compared with original data obtained during the initial waves of the SARS-CoV-2 pandemic.
Comprehending the factors influencing the in-host viral dynamics represents a fundamental tool to provide robust public health strategies. This pilot study could be implemented in further investigations involving other respiratory viruses, to better clarify the process of viral dynamics as a preparatory action for future pandemics.
2 Data and modeling assumptions
@@ 2.1 Samples and SARS-CoV-2 kinetics
Since February 2020, Fondazione IRCCS Policlinico San Matteo (Pavia, Italy) has been identified as reference regional center for the diagnosis of SARS-CoV-2 infections through the analysis of respiratory samples. In these analyses, RNA was extracted by using the MGISP-960 automated workstation and the MGI Easy Magnetic Beads Virus DNA/RNA Extraction Kit (MGI Technologies, Shenzhen, China). SARS-CoV-2 RNA was detected using the SARS-CoV-2 variants ELITe MGB kit (ELITechGroup, Puteaux, France; cat. no. RTS170ING) targeting ORF8 and RdRp gene. Reactions were carried out on the CFX96 Touch Real-time PCR detection system (BioRad, Mississauga, ON, Canada).
@@
$1\ \frac{\textrm{copies}}{\textrm{ml}}=16.6\times 10^{-0.2814\ \times\ 1\textrm{Ct}\ +\ 11.232},$ (1)
@@ 2.2 Statistical analysis
Numerical variables are described as mean $\pm$ standard deviation, categorical variables as count and percentages. The comparison of the two time periods (November–December 2020 and January–May 2021) considered is performed in terms of $t$-test for quantitative variables and Fisher exact test for quantitative variables. Considering the longitudinal structure of the data, they are presented as daily averaged values. The comparison between the two periods and the investigation of the influence of age and gender on the kinetics are performed in terms of a linear mixed-effects model. The temporal evolution of VL kinetics is obtained as a Gamma distribution using the Matlab built-in fit function, with the LAR method to ensure that the extreme values have a lesser influence on the result. The significance threshold is set to $p<0.05$.
@@ 2.3 Data description
2.4 Viral load kinetics
In this section we propose a reconstruction based on raw data of VL kinetics, to understand the main features of their evolution. Based on existing works in the medical literature about the shape of the SARS-CoV-2 infectiousness [13, 44] we expect the viral load to be characterized by a rapid growth in the few days after the first contact with an infective individual, until the infection reaches its maximum, followed by a slower decay up to the complete negativization of the subject after several days. We recall that a patient is considered negative if $\textrm{Ct}\geq 40$ is detected, based on the conversion formula (1). In order to compare VL kinetics from different patients, we determine the first day of negativization for each subject and we save all their preceding, patient-specific, VL measurements. Since there is no a priori information on the day when the peak of infectivity is reached, we align the different VL kinetics of all subjects based on the day of negativization, being this value determined within the whole dataset. In particular, the considered alignment does not depend on the observed peak of individual viral loads, which in turn is often the first measurement. Therefore, an alignment by negativization makes possible to observe an aggregate trend in the dataset. In Figure 1 we present the raw data of VL kinetics for the whole cohort of patients, aligned by the day of first negativization, over the two distinct time periods during which their infectiousness was monitored.
Figure 1: Evolution of individuals’ viral load kinetics over time with an ordering based on the first negative test taken by the infective patients, where a patient is considered negative if $\textrm{Ct}\geq 40$ is detected, based on the conversion formula (1). We use a simple linear interpolation to connect values that are relative to the same subject. We distinguish between the two observation periods of November–December 2020, accounting for 71 patients (left), and January–May 2021, accounting for 162 patients (right).
Based on this alignment process, we consider the daily average of the viral load with respect to the number of patients that got tested on the corresponding day. We then infer the evolution of the viral load as the best fit of these averages, weighted by the daily patients’ count. This means that the more patients got tested on one specific day, the more importance the corresponding average VL kinetics will have in the fitting. The outcomes of this analysis are shown in Figure 2. The infection persists for at most 25 days and its peak is reached 3 to 4 days after the first contact with an infected individual (4 days for the viral loads considered from the period November–December 2020 and 3 days for those belonging to the window January–May 2021). The red curves represent the best weighted fits for the daily averaged viral loads and have the shape of a Gamma distribution
@@ 2.4 Viral load kinetics
where $\alpha_{0}=10^{7}$ is a scale factor used to capture the magnitude of individuals’ viral load and the coefficients $\alpha_{1}\in\mathbb{R}$, $\alpha_{2}$, $\alpha_{3}\geq 0$ are reported in Table 1. In particular, $\alpha_{1}$ represents the time of first contact (measured in days), while $\alpha_{2}$ and $\alpha_{3}$ characterize the shape of the viral load, hence the probability of viral transmission between individuals.
@@
Table 1: Optimised coefficients for viral load kinetics, for the two observation periods of November–December 2020 (left) and January–May 2021 (right). The parameter $\alpha_{1}$ denotes the time of first contact (measured in days) while the shape parameters $\alpha_{2}$ and $\alpha_{3}$ characterize the probability of transmitting the infection.
@@ 2.5 Epidemic model with uncertain quantities
Compartmental models are classically defined to describe mathematically the spread of epidemics in a population [18]. The main idea of such models is to divide the population into different subgroups, each identified by a specific epidemiologically relevant status. Suitable transition rates are introduced to describe the switches between compartments. Generalizations of the mentioned compartmental models include additional factors characterizing the transmission dynamics such as external influence, age structure, variable contact dynamics and mobility of agents, see e.g. [1, 2, 11, 20]. From a model perspective, the choice of the compartmentalization is crucial to provide reliable predictions and, therefore, to effectively understand the evolution of the infection. In the following, we will adopt a SIR-type model with age of infection, where the compartment propagating the infection is the one identified by the population of infected agents. In particular, assuming the infectivity to be modulated by an additional variable identifying the age of infection, we can keep track of the disease’s variability inside each individual, as it evolves in parallel with the main timescale of the epidemic. In other words, we assume that the viral load heterogeneity is capable to characterize the population propagating the infection [28]. Experimental evidence suggests that in the growth phase of VL dynamics no symptoms are typically detected, see [38]. This allows to introduce a local incidence rate function of the age of infection, modeling the average viral load of the infected individuals and whose shape can be inferred by the information coming from real data on molecular tests. To take into account the structural uncertainties stemming from data assimilation processes of VL kinetics, we extend the model approach to consider the presence of uncertain quantities as a structural feature of the epidemic dynamics. To date, few results are available regarding the development of uncertainty quantification (UQ) methods in epidemic systems, see [1, 12, 15, 45]. The development of UQ methods are based on an increased dimensionality of the problem, that is made dependent on a random variable $\mathbf{z}\in\mathbb{R}^{d_{z}}$ of which the distribution $p(\mathbf{z})$ is known [27, 39, 48]. The extrapolation of statistics is classically obtained by looking at quantities of interest, with respect to $\mathbf{z}$, representing the expected solution of the problem.
Since the first day of infectivity of an infectious individual is unknown, we consider a compartmental modeling approach with uncertainty. We consider two temporal scales, one for the time $t\geq 0$ of evolution for the epidemic and another for the time $\tau\geq 0$ of evolution for the infection. At any given time $t\geq 0$, we then divide the population into susceptible individuals $S=S(t,\mathbf{z})$, recovered individuals $R=R(t,\mathbf{z})$ and infected individuals $i(t,\tau,\mathbf{z})$, with variable infection over time $\tau\geq 0$. We then consider the triple $(S(t,\mathbf{z}),i(t,\tau,\mathbf{z}),R(t,\mathbf{z}))$ of unknowns, solution to
@@
where $\beta(\tau,\mathbf{z})\geq 0$ and $\gamma(\tau,\mathbf{z})\geq 0$ represent respectively the uncertain infection rate and the uncertain recovery rate of the infected individuals, depending on the age of infection $\tau\geq 0$, while the constants $\bar{\beta}>0$ and $\bar{\gamma}>0$ enclose respectively the baseline contact rate and the baseline recovery rate. In (3) we indicated with $\partial_{t}$ the partial derivative with respect to the epidemic timescale $t\geq 0$ and with $\partial_{\tau}$ the partial derivative computed with respect to the infection timescale $\tau\geq 0$, see [17]. In particular, we stress once more that since the observable number of infected individuals also varies based on the age of infection $\tau\geq 0$, the quantity $i(t,\tau,\mathbf{z})\geq 0$ depends on both timescales (epidemic and infection course) and is related to the total current number of infected individuals $I=I(t,\mathbf{z})$ through the relation
@@
Therefore, this model couples the timescale of the epidemic with the one characterizing the evolution of the virus in terms of its average infectiousness. For each $\mathbf{z}$, the system is finally completed with suitable initial and boundary conditions for the variables $S(t,\mathbf{z})$, $i(t,\tau,\mathbf{z})$ and $R(t,\mathbf{z})$. A natural choice is
@@
and $p(\mathbf{z})$ is the distribution of the random variable $\mathbf{z}$.
The introduced simplified infection dynamics may be further modified to take into account a more complex compartmentalization of the society, the influence of external interventions and infection delays among others, see [10, 25, 24] and the references therein. In the following, we will always assume a constant $\gamma(\tau,\mathbf{z})\equiv 1$, so that the recovery rate is simply given by $\bar{\gamma}$. We point the interested reader to [51] for a more precise estimation of such parameters.
3 Inference of viral load dynamics
@@
In this section we define a pipeline to effectively estimate VL dynamics $\beta(\tau,\mathbf{z})$ by using available epidemiological data and introducing uncertainty in the day of first contact, expressed by the coefficient $\alpha_{1}$ in (2). To this end, we consider an uncertain parameter $\alpha_{1}(\mathbf{z})\in\mathbb{R}$ to make the local incidence rate random. Moreover, we take into account data from the COVID-19 pandemic in Italy, at both local and national level, by first considering the evolution of infected individuals in the province of Pavia and by then extending our analysis to the whole territory of Italy, where the central government implemented several non-pharmaceutical interventions (NPIs) during the pandemic. In Table 2 we report the timeline of the main NPIs taken before the mass vaccination campaign started in mid 2021. It is however worth mentioning that local additional measures, like isolation of small portions of a territory, had often been implemented as well.
@@
Before advancing to the estimation of VL dynamics, it is crucial to address some key points. We emphasize the inherent complexity in calibrating epidemiological models, which is particularly demanding from a numerical perspective. Indeed, the process of aligning these models with real-world scenarios requires a careful approach to parameter estimation. Furthermore, available data on VL kinetics tend to hide the initial evolution of their dynamics since agents are typically tested after the manifestation of the symptoms. This bias is primarily attributed to limited testing capacity, resulting in a dataset that often reflects a conservative estimate. Consequently, any attempts to determine the VL dynamics should be approached with a mindful consideration of these intrinsic limitations, coupled with a commitment to refining the model’s accuracy through continuous reassessment and adaptation.
Implemented NPI Date Average recovery time
First lockdown March 2020
Curfew measures October 2020 14 d
Strengthening of measures March 2021
@@
Table 2: Timeline of the main NPIs employed in Italy before the mass vaccination campaign started in mid 2021. The average recovery time from the infection was $14$ days.
@@
We shall focus on three different periods of time. We start by analyzing the first COVID-19 epidemic wave in Italy, which began in February–March 2020. We then proceed with the study of the second wave of October 2020 and conclude with the third major wave of February–March 2021. We stress that the data on VL kinetics are limited to the sole province of Pavia and have been collected during the two periods of November–December 2020 (characterized by the absence of VOCs) and January–May 2021 (involving the circulation of VOCs). Moreover, we point out that the epidemiological data at our disposal for the province of Pavia were only acquired up to January 18 2021 and we will therefore narrow down our analysis of the third epidemic wave solely to the case of Italy, for which we shall instead take advantage of the open access data collected by the Johns Hopkins University (github.com/CSSEGISandData/COVID-19), reporting in detail the daily number of infected, recovered and deceased individuals in Italy up to August 4 2021. Let $\mathcal{C}_{z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{M}\}$ be a sample of the random variable $\mathbf{z}$ having probability density function $p(\mathbf{z})$. We adopt a multi-level approach to estimate the parameters $(\alpha_{i})_{i=1,2,3}$ characterizing $\beta(\tau,\mathbf{z}_{j})\geq 0$, which consists of two steps.
@@
Step 1 – Determination of the epidemiological parameters. For starting, we initially assess the (deterministic) value of $\alpha_{1}$ dictated by the epidemic during the two weeks preceding any of the three main NPIs presented in Table 2 and we estimate at the same time the relevant epidemiological parameter $\bar{\beta}>0$ which defines the baseline contact rate, assuming no restrictions on social contacts were in place. Notice that we have to discard the variable $\mathbf{z}$ at this stage, because we first need to ensure that the day of first contact $\alpha_{1}$ is compatible with the epidemiological data, in order to quantify the shift between the infection and the epidemic timescales. For this, we look for the optimal $\alpha_{1}$ and $\bar{\beta}$ minimizing the relative weighted $L^{2}$ norm between the current number of infected and recovered individuals (in Pavia or in Italy, depending on the local or national setting under consideration), which we denote $\hat{I}(t)$ and $\hat{R}(t)$, and the theoretical evolution of these quantities, $I(t)=\int_{\mathbb{R}_{+}}i(\tau)\mathrm{d}\tau$ and $R(t)$, given by the model (3) in absence of uncertainty. We notice in particular that, since (3) only involves the recovered individuals inside the population, we take $\hat{R}(t)$ to be the sum of the reported numbers of recovered and deceased individuals. We thus solve the minimization problem
@@
over the time horizon $[t_{0},t_{f}]$ representing any of the three periods that can be identified with the initial spreading phase of the three epidemic waves under study: February 24–March 9 2020, October 7–October 22 2020 and February 24–March 10 2021. In particular, we choose here the infection rate $\beta(\tau)$ to be the Gamma function (2) obtained in the previous section from the viral loads measurements, rescaled by the factor $\alpha_{0}=10^{7}$. We use the fitting from November–December 2020 as initial guess for the optimization (6) during the first two epidemic waves, while we opt for the January–May 2021’s fitting to study the third wave. Notice that at this stage the parameters $\alpha_{2}$ and $\alpha_{3}$ are fixed from these two fits, since we only need to ensure a proper initial calibration of the time shift $\alpha_{1}$ to correctly align our model with the epidemiological data. Finally, we always consider a constant recovery rate $\bar{\gamma}=1/14$ to lower the complexity of the minimization. This value is consistent with the existing literature on the COVID-19 infection [14, 33, 51], for which the time to viral clearance has been reported to span approximately between 10 and 21 days on average over the different variants. In Table 3 we collect all the parameters obtained from this initial analysis, for each of the three waves.
@@
$\bar{\beta}$ 0.1486 $\textrm{d}^{-1}$ 0.1603 $\textrm{d}^{-1}$
$\bar{\gamma}$ 1/14 $\textrm{d}^{-1}$ 1/14 $\textrm{d}^{-1}$
@@
$\bar{\beta}$ 0.3992 $\textrm{d}^{-1}$ 0.1225 $\textrm{d}^{-1}$ 0.1323 $\textrm{d}^{-1}$
$\bar{\gamma}$ 1/14 $\textrm{d}^{-1}$ 1/14 $\textrm{d}^{-1}$ 1/14 $\textrm{d}^{-1}$
@@
Table 3: Epidemiological parameters and coefficients of the infectiousness function for the three periods of time under consideration, both at a local (province of Pavia) and a national level (Italy).